# 3-in-1: 2D Rotary Adaptation for Efficient Finetuning, Efficient Batching and Composability

### v1

@@ 1 Introduction
To assess the efficacy of RoAd, we perform comprehensive evaluations on the GLUE benchmark [53], eight commonsense reasoning tasks and four arithmetic reasoning tasks, utilizing RoBERTa [29] and LLaMA [49, 50] (Section §4.1). The results consistently show that RoAd surpasses other PEFT methods while maintaining a significantly reduced scale of trainable parameters ($<1\%$), as depicted in Figure 1. Additionally, RoAd employs element-wise rather than matrix multiplication, which notably improves throughput when serving heterogeneous requests within the same batch, achieving twice the throughput of LoRA [13] (Section §4.2). Furthermore, RoAd can be seamlessly integrated within an intervention framework [10], thereby enhancing model interpretability. We illustrate this through a composition experiment, demonstrating RoAd’s capacity to merge weights trained for different tasks and display a new capability (Section §4.3).
@@ 2.3 Intervention and composability
$\bm{b}$ denotes the hidden representation generated at row $i$ and column $k$ when the model processes an input, while $\bm{s}$ represents the corresponding representation when the model processes a different input. The matrix $\bm{R}\in\mathbb{R}^{r\times d_{1}}$, consisting of orthogonal rows, serves as a low-rank projection matrix where $d_{1}$ is the dimension of the representation and $r$ is the subspace dimension under intervention. Equation 1 illustrates the application of a DII to $\bm{b}$ using a counterfactual source representation $\bm{s}$ [^2] We adopt notation systems from Wu et al. [58].
@@ 3.1 Pilot study
$\bm{R}=\text{diag}(\bm{R}_{1},...,\bm{R}_{i},...,\bm{R}_{d_{2}/2})\quad\text{with}\quad\bm{R}_{i}=\begin{bmatrix}\cos\theta_{i}&-\sin\theta_{i}\\ \sin\theta_{i}&\cos\theta_{i}\\ \end{bmatrix}$ (2)
The trainable parameters are denoted as $\{\theta_{i}\}_{i=1}^{d_{2}/2}$. This 2D rotary adaptation involves rotating pairs of adjacent dimensions of $\bm{h}$, specifically dimensions $12i-$ 1 and $2i$, using the rotation matrix $\bm{R}_{i}$. [^5] The index in this work starts from 1 instead of 0. The rotation matrix $\bm{R}$ is characterized by its parameter efficiency, which is attributed to its sparse structure and the parameter sharing within each block $\bm{R}_{i}$. Additionally, $\bm{R}$ can be integrated directly into the existing pretrained weights, forming $\bm{W}=\bm{W}^{0}\bm{R}^{\top}$, which does not incur additional computational costs during inference. This design closely mirrors RoPE [48], with the notable difference that in our RoAd, $\theta_{i}$ is trainable and $\bm{R}_{i}$ does not incorporate positional information.
@@
We develop three variants of RoAd by altering the configuration of shared parameters as outlined in Table 1. RoAd1 introduces a minimal change to Equation 2 by incorporating a scaling factor $\alpha_{i}$. RoAd1 already shows impressive results for most tasks in Section §4.1. For some knowledge-intensive tasks, we observe that RoAd2 and RoAd4 obtain better results with more trainable parameters. To preserve the starting point of LLMs [22], we always initialize $\alpha_{i}=1$ and $\theta_{i}=0$.
Table 1: A summarization of three RoAd methods.
@@
Batching. In practice, we don’t need to save $\bm{R}$ as a sparse matrix and do matrix multiplication. Taking RoAd1 as an example, we only need to save two vectors: $\bm{R}^{1}=(\alpha_{1}\cos\theta_{1},\alpha_{1}\cos\theta_{1},\alpha_{2}\cos\theta_{2},\alpha_{2}\cos\theta_{2},...,\alpha_{d_{2}/2}\cos\theta_{d_{2}/2},\alpha_{d_{2}/2}\cos\theta_{d_{2}/2})$ and $\bm{R}^{2}=(\alpha_{1}\sin\theta_{1},\alpha_{1}\sin\theta_{1},\alpha_{2}\sin\theta_{2},\alpha_{2}\sin\theta_{2},...,\alpha_{d_{2}/2}\sin\theta_{d_{2}/2},\alpha_{d_{2}/2}\sin\theta_{d_{2}/2})$. Then $\bm{z}=\bm{Rh}=\bm{R}^{1}\otimes\bm{h}+\bm{R}^{2}\otimes\hat{\bm{h}}$, where $\hat{\bm{h}}$ is a rearranged version of $\bm{h}$ given by $\hat{\bm{h}}=(-h_{2},h_{1},-h_{4},h_{3},...,-h_{d_{2}},h_{d_{2}-1})$ and $\otimes$ denotes element-wise multiplication. This reformulation not only simplifies the representation of $\bm{R}$ but also enhances the efficiency of batching operations in RoAd, relying solely on element-wise multiplications rather than BMM.
Composability. RoAd can be incorporated into the DII framework as $\Phi(\bm{h})=\bm{Rh}=\bm{h}+\bm{R}(\bm{h}-\bm{R}^{\top}\bm{h})$, with $\bm{Rs}$ in Equation 1 being set to $\bm{h}$. Although a degree of relaxation is introduced to the orthogonality of $\bm{R}$, it is important to note that the rows of $\bm{R}$ remain orthogonal to each other within non-adjacent segments of the same block, $\bm{R}_{i}$. This offers a possibility for composability. We can finetune some rows on one task and other orthogonal rows on another task. Since they are orthogonal to each other, these two tasks should minimally affect each other, and the combination of these rows after finetuning could bring new multitasking learning ability.
@@ 4.1 Results on downstream tasks
In conclusion, for enhanced performance on downstream tasks and if a marginal increase in the storage capacity for trained parameters is acceptable, RoAd4 is the preferable option. Conversely, if the objective is to investigate how the model adjusts in terms of angle and magnitude, RoAd1 is recommended. Notably, all variants of RoAd incur the same computational overhead for batching.
@@ 4.2 Efficiency results for batching
We commence by highlighting the significance of weight merging for PEFT. Among the approaches discussed in Section §4.1, only LoRA [13], DoRA [25], BOFT [28], OFT [41], BitFit [59], (IA)3 [24], and our proposed RoAd enable the integration of trainable parameters with pretrained parameters without incurring additional inference overhead. As an illustration, we consider LoRA both with and without weight merging to underscore this process’s importance. Notably, the implementation of LoRA with merged weights effectively reverts to the original LLM. To assess throughput, we configure the system with a batch size of 1, generate 2048 tokens, and apply the LoRA modules across all linear layers. Figure 3 (Left) clearly illustrates that the unmerged LoRA exhibits a significantly smaller throughput compared to the merged LoRA. Additionally, it is evident that the throughput of the unmerged LoRA demonstrates only a weak correlation with the rank size, primarily due to the fact that the additional overhead is largely attributed to communication instead of computation.
@@
Figure 4: Qualitative comparison between RoAd and LoReFT for their composability. The prompt is always in English for different subspaces. Refer to Figure D.1, D.2 and D.3 for more examples.
@@ 4.3 Qualitative results for composability
In our investigation of RoAd’s ability to handle compositional tasks, we primarily engage in multilingual experiments similar to those conducted by Wu et al. [58]. We use two training datasets: a new version of HellaSwag [60] [^6] https://huggingface.co/datasets/LeoLM/HellaSwag_de, which comprises 1K samples with prompts in English and completions in German, and a 1K-sample subset of the Ultrafeedback [7] dataset, which focuses on instruction following tasks in English. Contrary to the above experiments that adapt the outputs of the linear layer, here we instead adapt the representations from the 16th block of LLaMA-7B, treating RoAd as a DII method. Specifically, we only adapt/intervene the representation of the final token in the prompt using RoAd1. We train the upper half of $\bm{R}$, i.e. $\{\bm{R}_{i}\}_{i=1}^{d_{2}/4}$, to handle the German completions in HellaSwag, and another half to complete the English sentences in Ultrafeedback. Both tasks are simultaneously trained but utilize distinct subspaces of $\bm{R}$. We train the model over five epochs with a learning rate of $35e-$ 3 and a batch size of 8.777The experiment is based on this notebook https://github.com/stanfordnlp/pyreft/blob/main/examples/composition/compreft.ipynb.
@@ Appendix A Limitations
We recognize that a primary constraint of our study is the restricted evaluation of downstream tasks, predominantly focused on natural language applications. This limitation primarily stems from constrained computational resources. In future work, we aim to expand our evaluation to include additional tasks such as image-to-text and text-to-image tasks. A second limitation pertains to the scalability of RoAd. Currently, it is not feasible to indefinitely increase the number of trainable parameters with RoAd. Nevertheless, our experiments demonstrate that RoAd4 already exhibits commendable performance. Exploring the integration of RoAd with other PEFT methods, such as LoRA, may offer potential avenues for enhancing scalability. We leave this to future work.

### v2

@@ 1 Introduction
To assess the efficacy of RoAd, we perform comprehensive evaluations on the GLUE benchmark [56], eight commonsense reasoning tasks and four arithmetic reasoning tasks, utilizing RoBERTa [31] and LLaMA [52, 53] (Section §4.1). The results consistently show that RoAd surpasses other PEFT methods while maintaining a significantly reduced scale of trainable parameters ($<0.1\%$), as depicted in Figure 1. Additionally, RoAd employs element-wise rather than matrix multiplication, which notably improves throughput when serving heterogeneous requests within the same batch, achieving twice the throughput of LoRA [14] (Section §4.2). Furthermore, RoAd can be seamlessly integrated within an intervention framework [11], thereby enhancing model interpretability. We illustrate this through a composition experiment, demonstrating RoAd’s capacity to merge weights trained for different tasks and display a new capability (Section §4.3).
@@ 2.3 Intervention and composability
$\bm{b}$ denotes the hidden representation generated at row $i$ and column $k$ when the model processes an input, while $\bm{s}$ represents the corresponding representation when the model processes a different input. The matrix $\bm{R}\in\mathbb{R}^{r\times d_{1}}$, consisting of orthogonal rows, serves as a low-rank projection matrix where $d_{1}$ is the dimension of the representation and $r$ is the subspace dimension under intervention. Equation (1) illustrates the application of a DII to $\bm{b}$ using a counterfactual source representation $\bm{s}$ [^2] We adopt notation systems from Wu et al. [61].
@@ 3.1 Pilot study
$\bm{R}=\text{diag}(\bm{R}_{1},\bm{R}_{2},...,\bm{R}_{d_{2}/2})\quad\text{with}\quad\bm{R}_{i}=\begin{bmatrix}\cos\theta_{i}&-\sin\theta_{i}\\ \sin\theta_{i}&\cos\theta_{i}\\ \end{bmatrix}$ (2)
The trainable parameters are denoted as $\{\theta_{i}\}_{i=1}^{d_{2}/2}$. This 2D rotary adaptation involves rotating pairs of adjacent dimensions of $\bm{h}$, specifically dimensions $12i-$ 1 and $2i$, using the rotation matrix $\bm{R}_{i}$. [^5] The index in this work starts from 1 instead of 0. The rotation matrix $\bm{R}$ is characterized by its parameter efficiency, which is attributed to its sparse structure and the parameter sharing within each block $\bm{R}_{i}$. Additionally, $\bm{R}$ can be integrated directly into the existing pretrained weights, forming $\bm{W}=\bm{W}^{0}\bm{R}^{\top}$, which does not incur additional computational costs during inference. This design closely mirrors RoPE [51], with the notable difference that in our RoAd, $\theta_{i}$ is trainable and $\bm{R}_{i}$ does not incorporate positional information. The overview of RoAd is shown in Figure 3.
@@
We develop three variants of RoAd by altering the configuration of shared parameters as outlined in Table 1. RoAd1 introduces a minimal change to Equation (2) by incorporating a scaling factor $\alpha_{i}$. RoAd1 already shows impressive results for most tasks in Section §4.1. For some knowledge-intensive tasks, we observe that RoAd2 and RoAd4 obtain better results with more trainable parameters. To preserve the starting point of LLMs [23], we always initialize $\alpha_{i}=1$ and $\theta_{i}=0$.
Table 1: A summarization of three RoAd variants.
@@
Batching. In practice, we don’t need to save $\bm{R}$ as a sparse matrix and do matrix multiplication. Taking RoAd1 as an example in Equation (4), we only save two vectors: $\bm{R}^{1}$ and $\bm{R}^{2}$. Then $\bm{z}=\bm{Rh}=\bm{R}^{1}\otimes\bm{h}+\bm{R}^{2}\otimes\hat{\bm{h}}$, where $\hat{\bm{h}}$ is a rearranged version of $\bm{h}$ and $\otimes$ denotes element-wise multiplication. This reformulation not only simplifies the representation of $\bm{R}$ but also enhances the efficiency of batching in RoAd, relying solely on element-wise multiplications rather than BMM.
$\bm{z}=\bm{Rh}$ $=\bm{R}^{1}\otimes\bm{h}+\bm{R}^{2}\otimes\hat{\bm{h}}$ (4)
$=\begin{bmatrix}\alpha_{1}\cos\theta_{1}\\ \alpha_{1}\cos\theta_{1}\\ \alpha_{2}\cos\theta_{2}\\ \alpha_{2}\cos\theta_{2}\\ \vdots\\ \alpha_{d_{2}/2}\cos\theta_{d_{2}/2}\\ \alpha_{d_{2}/2}\cos\theta_{d_{2}/2}\end{bmatrix}\otimes\begin{bmatrix}h_{1}\\ h_{2}\\ h_{3}\\ h_{4}\\ \vdots\\ h_{d_{2}-1}\\ h_{d_{2}}\end{bmatrix}+\begin{bmatrix}\alpha_{1}\sin\theta_{1}\\ \alpha_{1}\sin\theta_{1}\\ \alpha_{2}\sin\theta_{2}\\ \alpha_{2}\sin\theta_{2}\\ \vdots\\ \alpha_{d_{2}/2}\sin\theta_{d_{2}/2}\\ \alpha_{d_{2}/2}\sin\theta_{d_{2}/2}\end{bmatrix}\otimes\begin{bmatrix}-h_{2}\\ h_{1}\\ -h_{4}\\ h_{3}\\ \vdots\\ -h_{d_{2}}\\ h_{d_{2}-1}\end{bmatrix}$
Composability. RoAd can be incorporated into the DII framework as $\Phi(\bm{h})=\bm{Rh}=\bm{h}+\bm{R}(\bm{h}-\bm{R}^{\top}\bm{h})$, with $\bm{Rs}$ in Equation (1) being set to $\bm{h}$. Although a degree of relaxation is introduced to the orthogonality of $\bm{R}$, it is important to note that the rows of $\bm{R}$ remain orthogonal to each other within non-adjacent segments of the same block, $\bm{R}_{i}$. This offers a possibility for composability. We can finetune some rows on one task and other orthogonal rows on another task. Since they are orthogonal to each other, these two tasks should minimally affect each other, and the combination of these rows after finetuning could bring new multitasking learning ability.
@@ 4.1 Results on downstream tasks
Model Method #Paras. BoolQ PIQA SIQA HellaS. WinoG. ARC-e ARC-c OBQA Avg.
@@
Table 5: Score on AlpacaEval2.0 with LLaMA2-7B.
Method #Params. Finetuning Data Win Rate (%)
LoRA 0.83% 10K cleaned Alpaca 61.55
LoReFT 0.03% 10K cleaned Alpaca 60.21
RoAd1 0.02% 10K cleaned Alpaca 62.64
LoReFT 0.03% UltraFeedback [7] 61.68
RoAd1 0.02% UltraFeedback 62.60
Observed from the above-mentioned results, for enhanced performance on downstream tasks and if a marginal increase in the storage capacity for trained parameters is acceptable, RoAd4 is the preferable option. Conversely, if the objective is to investigate how the model adjusts in terms of angle and magnitude, RoAd1 is recommended. Notably, all variants of RoAd incur the same computational overhead for batching.
Instruction-following ability. We further benchmark RoAd using AlpacaEval2.0 [9]. We finetune LLaMA2-7B with two instruction-tuning datasets and evaluate the model using AlpacaEval2.0. This evaluation employs GPT-4 [39] to assess the responses generated by the finetuned model against those produced by Text-davinci-003. We don’t choose GPT-4 as the reference model, because GPT-4 is too powerful than LLaMA2-7B. The proof-of-concept experiment with LoRA shows the win-rate < 5%. As shown in Table 5, RoAd1 demonstrates superior performance compared to all baselines, while utilizing the least number of trainable parameters.
Multimodal ability. Lastly, we apply RoAd to the LLM backbone of LLaVA [26]. Liu et al. [26] requires 4.61% trainable parameters for LoRA on this task, while most tasks with LoRA in our paper need < 1%, showing that this task is knowledge-intensive. Therefore, we need to scale RoAd’s trainable parameters. For this purpose, we combine it with LoRA due to the limited number of $\theta_{i}$ and $\alpha_{i}$ in $\bm{R}$. The combination is represented as $\bm{z}=(\bm{R}\bm{W}^{0\top}+(\bm{BA})^{\top})\bm{x}$, where $\bm{A}$ and $\bm{B}$ are from LoRA. We adjust the LoRA rank to vary the number of trainable parameters. We combine RoAd1 with LoRA, but not RoAd2 or RoAd4, as their primary design purpose is to increase the number of trainable parameters.
Table 6: Visual instruction tuning results on LLaVA1.5-7B.
Method #Params. GQA SQA VQAT POPE Avg.
LoRA 4.61% 62.4 68.5 56.9 86.0 68.5
RoAd4 0.08% 60.0 66.9 53.3 85.5 66.4
RoAd1 + LoRA 1.19% 62.5 68.2 57.4 85.8 68.5
As shown in Table 6, with only 0.08% trainable parameters, RoAd4 already achieves 96.9% of the accuracy of LoRA with 4.61% trainable parameters. By combining RoAd1 with LoRA, we achieve the same performance as LoRA with only 1/4 of its trainable parameters. This demonstrates RoAd’s excellent scalability when combined with LoRA.
@@ 4.2 Efficiency results for batching
We commence by highlighting the significance of weight merging for PEFT. Among the approaches discussed in Section §4.1, only LoRA [14], DoRA [27], BOFT [30], OFT [44], BitFit [62], (IA)3 [25], and our proposed RoAd enable the integration of trainable parameters with pretrained parameters without incurring additional inference overhead. As an illustration, we consider LoRA both with and without weight merging to underscore this process’s importance. Notably, the implementation of LoRA with merged weights effectively reverts to the original LLM. To assess throughput, we configure the system with a batch size of 1, generate 2048 tokens, and apply the LoRA modules across all linear layers. Figure 4 (Left) clearly illustrates that the unmerged LoRA exhibits a significantly smaller throughput compared to the merged LoRA. Additionally, it is evident that the throughput of the unmerged LoRA demonstrates only a weak correlation with the rank size, primarily because the additional overhead is largely attributed to communication instead of computation.
@@
Figure 5: Qualitative comparison between RoAd and LoReFT for their composability. The prompt for different subspaces is always in English. Refer to Figure D.1, D.2 and D.3 for more examples.
@@ 4.3 Qualitative results for composability
In our investigation of RoAd’s ability to handle compositional tasks, we primarily engage in multilingual experiments similar to those conducted by Wu et al. [61]. We use two training datasets: a new version of HellaSwag [63] [^6] https://huggingface.co/datasets/LeoLM/HellaSwag_de, which comprises 1K samples with prompts in English and completions in German, and a 1K-sample subset of the Ultrafeedback [7] dataset, which focuses on instruction following tasks in English. Contrary to the above experiments that adapt the outputs of the linear layer, here we instead adopt the representations from the 16th block of LLaMA-7B, treating RoAd as a DII method. Specifically, we only adapt/intervene the representation of the final token in the prompt using RoAd1. We train the upper half of $\bm{R}$, i.e. $\{\bm{R}_{i}\}_{i=1}^{d_{2}/4}$, to handle the German completions in HellaSwag, and another half to complete the English sentences in Ultrafeedback. Both tasks are simultaneously trained but utilize distinct subspaces of $\bm{R}$. We train the model over five epochs with a learning rate of $35e-$ 3 and a batch size of 8.777The experiment is based on this notebook https://github.com/stanfordnlp/pyreft/blob/main/examples/composition/compreft.ipynb.
@@ 5 Conclusion
Acknowledgements
We thank eBay Inc. for the computation support. This research was funded in part by the Netherlands Organization for Scientific Research (NWO) under project number VI.C.192.080.
@@ Appendix A Limitations
We recognize that a primary limitation pertains to the scalability of RoAd. Currently, it is not feasible to indefinitely increase the number of trainable parameters with RoAd. Nevertheless, our experiments demonstrate that RoAd4 already exhibits commendable performance. To scale the trainable parameters, we can combine RoAd with other PEFT methods, such as LoRA, which enhances the scaling behavior of these PEFTs, i.e. achieving similar results with less trainable parameters.